OpenStudy (anonymous):

solve by completing the square. 8x^2+22x=21

OpenStudy (anonymous):

divide by 8 to get \[x^2+\frac{11}{4}x=\frac{21}{8}\] then \[(x+\frac{11}{8})^2=\frac{21}{8}+\frac{121}{64}\] \[(x+\frac{11}{8})^2=\frac{289}{64}\] \[x+\frac{11}{8}=\pm\sqrt{\frac{289}{64}}\] \[x+\frac{11}{8}=\pm\frac{17}{8}\]

OpenStudy (anonymous):

and so \[x=\frac{17}{8}-\frac{11}{8}=\frac{6}{8}=\frac{3}{4}\] or \[x=\frac{-17}{8}-\frac{11}{8}=\frac{-28}{8}=-\frac{7}{2}\]